OpenStudy (anonymous):

It takes 3/7 of a carton of milk to fill a large glass. how many glasses can be filled using 50 cartons of milk? Show all Working out???

OpenStudy (anikhalder):

3/7 of a carton fills 1 glass So, I carton fills=\[1\div \frac{ 3 }{ 7}=7/3 \]glasses Hence, 50 such cartons will fill=\[50\times \frac{ 7 }{ 3}=116\frac{ 2 }{ 3}\]glasses. So we can say that 50 cartons will fill 116 glasses and 2/3 of the 117th glass.